OpenStudy (anonymous):

Help solving this identity? Medal and fan 2sin(2x)-2sinx +2sqrt3cosx-sqrt3=0 Thanks(:

mathslover (mathslover):

sin(2x) = ?

Miracrown (miracrown):

equation Well, it may be an identity, but I don't think so.

Miracrown (miracrown):

If it were an identity it would be true for all values of x

mathslover (mathslover):

Just wrongly phrased.

hartnn (hartnn):

the fact that left side expression can be easily factored, makes this quite easy to solve. ofcourse after applying the formula for sin 2x :)

Miracrown (miracrown):

So the first thing I would do is turn sin(2x) into 2 sin(x)cos(x) 2 (2 sin(x) cos(x) ) - 2 sin(x) + 2 Sqrt[3] cos(x) - Sqrt[3] = 0. 4 sin(x) cos(x) - 2 sin(x) + 2 Sqrt[3] cos(x) - Sqrt[3] = 0.

mathslover (mathslover):

Mira, let the user do it herself! I don't find any reason to put all of our work unless the user makes an effort.

OpenStudy (anonymous):

Sorry, I was afk. Yeah, I didn't mean to put identity, It's 3:24 am, a bit tired. Right, I have already done that in my work.

mathslover (mathslover):

Okay, so, is that what you have done yet? or something else too?

OpenStudy (anonymous):

Yep, I have that.

OpenStudy (anonymous):

I then took a 2sinx out...would that be right to do?

OpenStudy (anonymous):

Oh wait, nvm, I shouldn't do that.

OpenStudy (anonymous):

What direction should I be going with this? I need a push forward, I already had all the information that people had given.

mathslover (mathslover):

Let us start from the beginning. What is the formula for sin(2x) = ?

OpenStudy (anonymous):

sin(2x) = 2sinxcosx

mathslover (mathslover):

Good. Now, put this in the equation we have.

OpenStudy (anonymous):

(4sinxcosx-2sinx+ 2sqrt3cosx-sqrt3 = 0

mathslover (mathslover):

Good. now, let us try to factor it out.. the last two terms have something common in them, what is it?

OpenStudy (anonymous):

the sqrt3.

mathslover (mathslover):

Good work! Now, in the first two terms, what is common?

OpenStudy (anonymous):

the 2sinx

mathslover (mathslover):

Good work again...! Now, try to do the factoring. Can you do it?

OpenStudy (anonymous):

(sqrt3)(2sinx)(2sinxcosx+2cosx-2) so far correct?

mathslover (mathslover):

No! This is not the correct way to do the factoring. See, we have 2sin x common int he first 2 terms and sqrt(3) common in the last 2 terms : (4sinxcosx-2sinx+ 2sqrt3cosx-sqrt3 = 0 \(2\sin x ( 2\cos x - 1) + \sqrt{3} (2\cos x - 1) =0 \) Are you getting it?

OpenStudy (anonymous):

Yeah, I get that sorry. Factoring with trig functions always throws me off.

mathslover (mathslover):

No worries! So, just take 2cos(x) - 1 common now.

OpenStudy (anonymous):

Hold on, how did you get rid of the other 2sinx attached to the cos when it was 4sinx?

mathslover (mathslover):

\(4\sin x \cos x - \sin x + 2\sqrt{3} \cos x - \sqrt{3} = 0\) See, the first two terms have 2 sin x common : So, if you take 2 sin x common from 4 sin x cos x you get : 2 cos x

mathslover (mathslover):

Thus, I wrote it as: 2sinx (2cos -1) + sqrt(3) (2 cos x-1)= 0

OpenStudy (anonymous):

That still doesn't account for the 4 though..taking 2sinx out leaves 2sinxcosx...unless you converted the 2sinx into 2cosx-1, but then you also have the cosx still left there to multiply?

OpenStudy (anonymous):

(2sinxcosx) = 2sinx ((2cosx-1)cosx-1) ?

mathslover (mathslover):

Let us take only 2 terms now : 4 sin x cos x - sinx Now, I can write 4sinx cosx as : \(2\sin x \times 2\cos x \) Right?

OpenStudy (anonymous):

OH. I was only connecting that with the sin only. That makes much more sense.

mathslover (mathslover):

Oops wait, I wrote that wrng it is 4sinx cos x - 2sinx

mathslover (mathslover):

Now, we have : \(\color{blue}{2 \sin x} \times 2\cos x - \color{blue}{2\sin x}\) Take 2sinx common now..

OpenStudy (anonymous):

Right.

mathslover (mathslover):

So, what do you get?

OpenStudy (anonymous):

Well, I'm guessing we don't pay any mind to the sqrt 3, so I have 2sinx and cosx= 1/2

mathslover (mathslover):

No...! Don't jump directly to the conclusion, as this may lead to big mistakes. We have : \(2\sin x ( 2\cos x - 1) + \sqrt{3} (2\cos x -1 )=0\) Take \(2\cos x - 1\) common.

Miracrown (miracrown):

So the first thing I would do is turn sin(2x) into 2 sin(x)cos(x) 2 (2 sin(x) cos(x) ) - 2 sin(x) + 2 Sqrt[3] cos(x) - Sqrt[3] = 0. 4 sin(x) cos(x) - 2 sin(x) + 2 Sqrt[3] cos(x) - Sqrt[3] = 0. then I would factor by grouping. 2 sin(x) (2 cos(x) - 1) + Sqrt[3] (2 cos(x) - 1) = 0 Therefore (2 sin(x) + sqrt[3])(2 cos(x) - 1) = 0. This then is easily solvable for the roots. cos(x) = 1/2 So this means +- Pi/3 + 2 * n * Pi where n is any integer. sin(x) = - Sqrt[3]/2. This means 4pi/3 + 2 * n * Pi, n any integer. We ignore the other solution 5pi/3 + 2 * n * Pi because it's redundant with what was found from the other root.

Miracrown (miracrown):

Make sense? :) @seehearcreate

OpenStudy (anonymous):

Oh, you multiply the sinx by sqrt 3..... Yeah, that makes sense(:

Miracrown (miracrown):

yup! =)

mathslover (mathslover):

Sorry, but I don't appreciate someone jumping in and writing the whole direct solution. Anyways, no worries, I will leave. Good luck @seehearcreate !

OpenStudy (anonymous):

@mathslover You helped me the most with this problem, so I'll medal you, but I'll fan you both(: Thankyou, both of you!

Miracrown (miracrown):

You're welcome! =)